Theory and Applications of Categories, Vol. 17, No. 1, 2006, pp. 10–16.

THE CHU CONSTRUCTION: HISTORY OF AN IDEA

MICHAEL BARR

Abstract. This paper describes the historical background and motivation involved in the discovery (or invention) of Chu categories.

In 1975, I began a sabbatical leave at the ETH in Zürich, with the idea of studying duality in categories in some depth. By this, I meant not such things as the duality between Boolean algebras and Stone spaces, nor between compact and discrete abelian groups, but rather self-dual categories such as complete semi-lattices, finite abelian groups, and locally compact abelian groups. Moreover, I was interested in the possibilities of having a category that was not only self dual but one that had an internal hom and for which the duality was implemented as the internal hom into a “dualizing object”. This was already true for the complete semi-lattices, but not for finite abelian groups or locally compact abelian groups. The category of finite abelian groups has an internal hom, but lacks a dualizing object, while locally compact groups have a dualizing object, but not an internal hom that is defined everywhere. Although you could define an abelian group of continuous homomorphisms between locally compact abelian groups, there was no way of systematically putting a locally compact topology on the hom set that would lead to the desired properties.

The desired properties were what I subsequently called ∗-autonomy. I do not want to go into the technical details here (see [Barr, 1999], for example). Basically, a ∗-autonomous category has an internal hom, denoted −∘, a symmetric monoidal structure, ⊗, along with the usual coherent isomorphisms Hom(A ⊗ B, C) ≃ Hom(A, B −∘ C). In addition, there should be a “dualizing object”, ⊥, such that if we define A^{∗} = A −∘ ⊥, the subsequent canonical map (A −∘ B) [figure] (B^{∗} −∘ A^{∗}) should be an isomorphism.

There is a certain redundancy in this definition. For instance, the most efficient (and perhaps the most natural) way is to assume just −∘ and ⊥ and define A^{∗} = A −∘ ⊥ and then A ⊗ B = (A −∘ B^{∗})^{∗}, since the latter two are provably naturally isomorphic.

By the end of the year, I had in fact produced a moderate number of examples of ∗- autonomous categories. One of them was a full subcategory of topological abelian groups that included all the locally compact abelian (LCA) groups in such a way that the duality restricted to them was the well-known duality of LCA groups. As one would expect, the the circle group was the dualizing object. In addition, the category was complete and cocomplete. In fact, it was built, essentially, by completing the LCA groups.

This work was supported by a grant from the NSERC of Canada

Received by the editors 2003-06-06 and, in revised form, 2006-11-09.

Published on 2006-12-16 in the volume Chu spaces: theory and applications.

2000 Mathematics Subject Classification: 18-03, 01A65, 18D15.

Key words and phrases: Chu categories, *-autonomous categories, history.

\begin{matrix}c \\ ◯ \end{matrix} Michael Barr, 2006. Permission to copy for private use granted.

Another example was a full subcategory of the category of locally convex topological vector spaces. In the process of studying that category, I read a number of books on that subject, for example [Schaefer, 1970] and [Pietsch, 1972]. Each one contained a certain construction of “pairs”. A pair (E, E′) consists of two (discrete) vector spaces equipped with a bilinear map^{1} ⟨−, −⟩ : E × E′ [figure] C, the complex numbers. This device was used because of the large number of possible topologies on the dual space of a topological vector space. Basically, you were to think of E and E′ as being dual to each other with certain topologies.

As I was preparing an early draft of the subsequent monograph [Barr, 1979], I began to be concerned about the fact that I had constructed this elaborate theory and had only a bare handful of examples. Aside from the complete semi-lattices the other examples were constructed with some difficulty using quite ad hoc methods. I had noticed the pairs and wondered if there were a way of defining an internal hom among pairs. It turned out to be surprisingly easy. I began with the obvious interpretation that in a pair E = (E, E′), E was the set of points and E′ the set of functionals. (Of course, in the dual pair E′ is the set of points and E the set of functionals.) So if F = (F, F′) is another such object, the points of E −∘ F should certainly be the set of maps between the pairs and a map was a continuous linear map f : E [figure] F such that for all ϕ ∈ F′, ϕ.f ∈ E′. Call this set [E, F]. Then we wanted

E −∘ F = ([E, F],?)  (∗)

with ? an unknown to be determined. A similar guess was that the tensor product of two objects might be given by

(E, E′) ⊗ (F, F′) = (E ⊗ F,??)  (∗∗)

where ?? is another unknown to be determined. But if we put these together with the fact that in any ∗-autonomous category,

E ⊗ F = (E −∘ F^{∗})^{∗}

(or, equivalently, that E −∘ F = (E ⊗ F^{∗})^{∗}), we can compute

E ⊗ F = (E −∘ F^{∗})^{∗}
= ((E, E′) −∘ (F′, F))^{∗}
= ([(E, E′), (F′, F)],?)^{∗}
= (?, [(E, E′), (F′, F)])

Comparing this with (∗∗) we see that

E ⊗ F = (E ⊗ F, [(E, E′), (F′, F)])

1It was usually supposed that the pairing was non-singular. As a first approximation, I had to omit that hypothesis. Later, a way was found to reinstate it, if desired, at least in the familiar examples.

Similarly, we find that

E −∘ F = ([E, F], E ⊗ F′)

These formulas are suggestive, but they don’t prove anything. But, amazingly, it all works!

Thus I ended up with a category whose objects were pairs E = (E, E′) of vector spaces equipped with a pairing E ⊗ E′ [figure] C. A map from E to F = (F, F′) is a pair of linear maps (f, f′) in which f : E [figure] F and f′ : F′ [figure] E′ (note the direction reversal) such that ⟨fv, w⟩ = ⟨v, f′w⟩ whenever v ∈ E and w ∈ F′. There is no topology assumed and no continuity on the linear maps.

At this point, it seemed clear that no special properties of vector spaces was being used and this construction was an instance of a general one. Let 𝒱 be an autonomous category, meaning one which had a coherently associative, unitary, symmetric tensor product, − ⊗ −, an internal hom − −∘ − with the usual adjunction

Hom(A ⊗ B, C) ≃ Hom(A, B −∘ C)

The only additional assumption we need to make on 𝒱 is that it have pullbacks. Fix an object of 𝒱 , which we will call ⊥. Form the category of pairs E = (E, E′) equipped with

a morphism E ⊗ E′  [figure] ⊥. If F = (F, F′) is another object, then a morphism E  [figure] F
is a pair (f, f′) where f : E  [figure] F and f′ : F′  [figure] E′ such that the diagram

[figure: E ⊗ E′; ⊥; ⟨−,−⟩; E ⊗ F′; E ⊗ E′; E⊗f′; E ⊗ F′; F ⊗ F′; f⊗F′; F ⊗ F′; ⊥; ⟨−,−⟩]

commutes. This condition is just a diagrammatic interpretation of the definition of morphism. It is equivalent to the assertion that the following square is a pullback:

[figure: Hom(F′, E′); Hom(E ⊗ F′, ⊥); Hom(E, F); Hom(F′, E′); Hom(E, F); Hom(E, F); Hom(E, F); Hom(E ⊗ F′, ⊥)]

This definition gives a category that is called Chu(𝒱 , ⊥). Assuming that 𝒱 has pullbacks, we can internalize the above definition, by defining [E, F] so that

[figure: F′ −∘ E′; E ⊗ F′ −∘ ⊥; [E, F]; F′ −∘ E′; [E, F]; E −∘ F; E −∘ F; E ⊗ F′ −∘ ⊥]

is a pullback. Then define

E −∘ F = ([E, F], E ⊗ F′)

The pairing is either of the two equal maps

[E, F] ⊗ E ⊗ F′  [figure] (E −∘ F) ⊗ E ⊗ F′  [figure] F ⊗ F′  [figure] ⊥

or

[E, F] ⊗ E ⊗ F′  [figure] (F′ −∘ E′) ⊗ E ⊗ F′  [figure] E′ ⊗ E  [figure] ⊥

It seemed clear that this gave a ∗-autonomous category, but there were a number of unpleasant details to be verified. Since my student, Po-Hsaing Chu needed a master’s project, so I asked him to verify them, which he did [Chu, 1978, 1979]. I now had expanded from six to infinity the repertory of known ∗-autonomous categories. At this point, the construction, and in fact the whole idea of ∗-autonomous categories, was more-or-less forgotten, not least by me.

One interesting thing is that I never looked to see if there was any connection between the ∗-autonomous categories that I had carefully constructed and related ∗-autonomous categories. For example, I had constructed a ∗-autonomous category that was a full subcategory of topological abelian groups and made no attempt to relate it to Chu(Ab, K), where K is the circle group. This was a mistake.

Eight years later, in 1987, I became aware of Jean-Yves Girard’s linear logic. I do not recall how I became aware of it, but at the meeting at The University of Colorado that summer, I knew enough to bring Girard a copy of [Barr, 1979]. Since ∗-autonomous categories give models of linear logic and since the Chu construction gives easy examples of ∗-autonomous categories, this was reasonable, but I must have known that before going to Boulder. I think Robert Seely must have lectured on this in our seminar that spring. At any rate, knowledge of ∗-autonomous categories in general and the Chu construction in particular had penetrated the linear logic and theoretical computer science community.

1. Further work

Vaughan Pratt is convinced that Chu categories are an important adjunct to the study of a number of questions in theoretical computer science. He has written a series of papers expanding on this perception. A small selection of them is [Pratt, 1993a, 1993b, 1995].

In the meantime, I became interested in the connection between the categories constructed in my 1979 monograph and the Chu construction. In that paper, I had constructed a small number of ∗-autonomous categories. In all but one case I had begun with an incomplete, but ∗-autonomous category that was embedded as a full subcategory of a complete autonomous category that lacked a duality. I then showed how there was a kind of completion of the ∗-autonomous subcategory inside the category of topological (or uniform space) objects of the complete autonomous category that was ∗-autonomous. The construction was fairly difficult and it was not easy to understand which objects are in the subcategory. Eventually, it occurred to me that if 𝒱 is the category of discrete objects and ⊥ is the dualizing object, then Chu(𝒱 , ⊥) is a candidate for the ∗-autonomous category that I had constructed.

For example, in the category of abelian groups, there is a full ∗-autonomous subcategory of the category of topological abelian groups that includes all the locally compact groups. Let chu(Ab, K) denote the full subcategory of Chu(Ab, K) consisting of those pairs (A, A′) for which the pairing is non-singular: for any a ∈ A, if a ≠ 0, then there is an a′ ∈ A′ for which ⟨a, a′⟩ ≠ 0 and vice versa. The definitions of the internal hom and tensor have to be modified, but the modifications are essentially obvious. It is now known that chu(Ab, K) is equivalent to a full subcategory of topological abelian groups that includes all the locally compact group. Is this the same category as constructed “by hand”? I don’t know and I have no idea how to even approach the question (and, to be sure, there is not any reason to know the answer).

2. Mackey

I had imagined that the source of the idea of pairs of vector spaces was Grothendieck. It looked like the kind of construction he would have done, although he might have made a category of it. But in fact, it goes back further than that. The source seems to have been George Mackey’s thesis, published as [Mackey, 1945]. In it he defines pairs as above, except that his pairs are not assumed to be non-singular in the second variable. So the dual of (E, F) is (F, E/E_{0}) where E_{0} consists of those vectors that are annihilated of every functional in F.

One interesting point about Mackey’s paper is that it looks as if he was proposing that the set of functionals be considered as a replacement for the topology, not just an adjunct to it. After all, the topology on a set is a equivalent to a set of functions from the set into the Sierpinski space 2. Not any set of functions will do; it must be a ⋃-∩-closed sublattice of the set of all functions. In the case of the vector spaces, the condition is even simpler; it must be a linear subspace of the space of all functionals. Note that I am not suggesting that this is the same thing as a topological vector space. It is not, but it might be a subject of more interest, properly studied on its own, than topological vector spaces. At all events, if this was Mackey’s intention, it has not been followed up^{2}. Instead, the various authors have used these pairs as a tool for the study of duality, rather than as the objects of study. And the dual is, in turn, used as a tool in the study of the deeper properties of these spaces.

3. A rational—but incorrect—reconstruction

Here is how the Chu construction should have been discovered. That it wasn’t just shows how sometimes mathematics is two steps forward, then one back.

2The referee informs me that some people have and mentioned books by Frölicher and Kriegl, and of Kriegl and Michor. Interestingly Frölicher heard me lecture in this topic and didn’t mention the book.

Suppose you want to make a self-dual category that contains a given category 𝒞. Assuming that 𝒞 has a terminal object 1, the category 𝒞^{e} = 𝒞 × 𝒞^{op} is obviously self-dual and embeds 𝒞 as the objects of the form (C, 1). If 𝒞 should happen to be (symmetric) closed monoidal, then you might not expect 𝒞^{e} to be a closed monoidal category, but it is, at least when 𝒞 has finite products. The reason you might not expect it is that 𝒞^{op} is not generally closed monoidal when 𝒞 is. For instance, the category of finite dimensional coalgebras over a field is closed monoidal (in fact, cartesian closed), but the opposite category of algebras certainly is not. What is true (and it is crucial) is that when 𝒞 is closed monoidal, 𝒞^{op} is enriched over 𝒞.

We begin with

Hom((C, D), (C′, D′)) = Hom(C, C′) × Hom(D′, D)

which is the definition of the homfunctor in 𝒞^{e}. This leads one to try

(C, D) −∘ (C′, D′) = ((C −∘ C′) × (D′ −∘ D), X)  (∗)

where X = X(C, D, C′, D′) is to be determined. Similarly, you might guess that

(C, D) ⊗ (C′, D′) = (C ⊗ C′, Y)

where Y = Y (C, D, C′, D′) is to be determined. Naturally, you want more than just self dual and closed monoidal, you want ∗-autonomous, which implies that

\begin{aligned}(C, D) −∘ (C′, D′) & ≃ ((C, D) ⊗ (C′, D′)^{∗})^{∗} = ((C, D) ⊗ (D′, C′))^{∗} \\ & = ((C, D) ⊗ (D′, C′), Y (C, D, D′, C′))^{∗} = (Y (C, D, D′, C′), C ⊗ D′) \end{aligned}

Comparing this to (∗), we see that X(C, D, C′, D′) = C ⊗ D′ and Y (C, D, D′, C′) = (C −∘ C′) × (D′ −∘ D). This leads directly to the formulas

(C, D) −∘ (C′, D′) = ((C −∘ C′) × (D′ −∘ D), C ⊗ D′)

and

(C, D) ⊗ (C′, D′) = (C ⊗ C′, (C −∘ D′) × (C′ −∘ D))

Amazingly, this all works! In fact, the reader might have noticed that it is just Chu(𝒞 , 1); an object of the latter category is a pair (C, D), together with a morphism C ⊗ D [figure] 1, of which there is exactly one. This category could thus have been discovered with no motivation from Mackey’s pairs. Could have been but wasn’t.

Moreover, this leads directly to the Chu construction. If M is a commutative monoid object in any ∗-autonomous category, an M-action is an object A, together with an morphism M ⊗ A [figure] A that satisfying the usual associative and unitary identities. It is not hard to prove that the category of M-actions is again a ∗-autonomous category. Any object ⊥ of 𝒞 gives a monoid object M = (⊤, ⊥) of 𝒞^{e}. Since the unit for the tensor

16  MICHAEL BARR

product of 𝒞^{e} is (⊤, 1) and the unit map for M is (id,!), where ! : ⊥ [figure] 1 is the unique arrow. The multiplication is

(⊤, ⊥) ⊗ (⊤, ⊥) = (⊤ ⊗ ⊤, (⊤ −∘ ⊥) × (⊤ −∘ ⊥)) = (⊤, ⊥ × ⊥) [figure: (id,Δ)] (⊤, ⊥)

It then turns out that the category of M-actions is precisely Chu(𝒞 , ⊥).

Moreover, this constructions works with certain changes even when ⊗ is not symmetric. In that case, two internal homs are needed in 𝒞, one to be right adjoint to C ⊗ − and the other to be right adjoint to − ⊗ C. In that case, M need not be commutative (in fact, the concept is undefinable) and instead of left actions of M, we must use the category of two-sided actions. This is studied in detail in [Barr, 1995], where the ∗-autonomous structure of 𝒞^{e} was first described. The simplification that this construction offers is more important in the asymmetric case than in the symmetric one.

References

M. Barr (1979), ∗-Autonomous categories. Lecture Notes in Mathematics 752 Springer-Verlag, Berlin,

Heidelberg, New York. M. Barr (1995), Non-symmetric ∗-autonomous categories. Theoretical Computer Science, 139, 115-130. M. Barr (1996), The Chu construction. Theory and Applications of Categories, 2, 17–35. M. Barr (1999), ∗-Autonomous categories: once more around the track. Theory and Applications of

Categories, 6, 5–24. M. Barr (2000), On ∗-autonomous categories of topological vector spaces. Cahiers de Topologie et Géo-

métrie Différentielle Catégorique, 39 243–254. P.-H. Chu (1978), Constructing ∗-autonomous categories. M. Sc. thesis, McGill University. P.-H. Chu (1979), Constructing ∗-autonomous categories. Appendix to [Barr, 1979]. G. Mackey (1945), On infinite dimensional vector spaces. Trans. Amer. Math. Soc. 57, 155–207. A. Pietsch, Nuclear Locally Convex Spaces. Springer-Verlag, 1972. V.R. Pratt (1993a), Linear Logic for Generalized Quantum Mechanics. In Proceedings Workshop on

Physics and Computation, Dallas, IEEE Computer Society. V.R. Pratt (1993b), The Second Calculus of Binary Relations. In Proceedings of MFCS’93, Gdańsk,

Poland, 142–155. V.R. Pratt (1995), The Stone Gamut: A Coordinatization of Mathematics. In Proceedings of the confer-

ence Logic in Computer Science IEEE Computer Society. H.H. Schaefer, Topological Vector Spaces. Third printing, Springer-Verlag, 1970.

Department of mathematics and statistics, McGill University 805 Sherbrooke St. W, Montreal, QC, H3P 1S4

Email: barr@barrs.org

This article may be accessed via WWW at http://www.tac.mta.ca/tac/ or by anonymous ftp at ftp://

ftp.tac.mta.ca/pub/tac/html/volumes/17/1/17-01.{dvi,ps,pdf}

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. The method of distribution of the journal is via the Internet tools WWW/ftp. The journal is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive (by e-mail) abstracts of articles as they are published. Full text of published articles is available in .dvi, Postscript and PDF. Details will be e-mailed to new subscribers. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh,

rrosebrugh@mta.ca.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX2e is the preferred flavour. TEX source of articles for publication should be submitted by e-mail directly to an appropriate Editor. They are listed below. Please obtain detailed information on submission format and style files from the journal’s WWW server at http://www.tac.mta.ca/tac/. You may also write to tac@mta.ca to receive details by e-mail.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: mbarr@barrs.org

Transmitting editors.

Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: r.brown@bangor.ac.uk Aurelio Carboni, Università dell Insubria: aurelio.carboni@uninsubria.it Valeria de Paiva, Xerox Palo Alto Research Center: paiva@parc.xerox.com

Ezra Getzler, Northwestern University: getzler(at)math(dot)northwestern(dot)edu

Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk P. T. Johnstone, University of Cambridge: ptj@dpmms.cam.ac.uk G. Max Kelly, University of Sydney: maxk@maths.usyd.edu.au Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, University of Western Sydney: s.lack@uws.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@acsu.buffalo.edu Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Brooke Shipley, University of Illinois at Chicago: bshipley@math.uic.edu James Stasheff, University of North Carolina: jds@math.unc.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: robert.walters@uninsubria.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca